What is another equation that would be parallel to the equation y=x-3 and pass through the points (2,1) ?
Answer by stanbon(75887)

What is another equation that would be parallel to the equation y=x-3 and pass through the points (2,1) ?
---
The slope of the given equation is 1.
---
You want an equation with slope = 1 where y = 1 when x = 2:
y = mx + b
Solve for "b":
1 = 1*2 + b
b = -1
-----
Equation:
y = x -1
===============
Cheers,
Stan H.
===============